OpenStudy (anonymous):

Find the digit that makes 3,71__ divisible by 9. A)3 B)7 C)1 D)5

OpenStudy (anonymous):

just add it at the end

OpenStudy (anonymous):

yeah what digit

OpenStudy (anonymous):

7

OpenStudy (anonymous):

use a calculator to divide it by 9

OpenStudy (anonymous):

ok i have another question

OpenStudy (anonymous):

it goes into it evenly

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Find the prime factorization of the number 168

OpenStudy (anonymous):

is that the 1 with the factor tree

OpenStudy (anonymous):

2*2*2*3*7

OpenStudy (anonymous):

2^3*7*3

OpenStudy (anonymous):

whichever 1 they both work

OpenStudy (anonymous):

here are the answers: a. 2 to the 4th power x 3x7 b.2 to the 3rd power x 3 x 7 c.2 to the 4th power x 3 to the 3 power x 13 d.2 to the 3rd power x 3 to the 3rd power x 7

OpenStudy (anonymous):

b

OpenStudy (anonymous):

so what awnser is it

OpenStudy (anonymous):

thats the same thing as 2^3*7*3

OpenStudy (anonymous):

so b

OpenStudy (anonymous):

3717/9 = 413 3+7+1+7=18=9*2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok got another

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

find the GCF of the numbers 140 and 180 a.20 b.90 c.30 d.10

OpenStudy (anonymous):

20 is the biggest number that can go into them both

OpenStudy (anonymous):

anything else

OpenStudy (anonymous):

identify the fraction that is equivalent to the given fraction. 4/7 a. 16/28 b.20/28 c.16/21 d.12/28

OpenStudy (anonymous):

4*4=16 7*4=28

OpenStudy (anonymous):

16/28

OpenStudy (anonymous):

anymore?

OpenStudy (anonymous):

ernest can walk a distance of 5 miles in 75 minutes. which is the simplest form of 5/75? a.75/5 b.2/25 c.1/15 d.1/70

OpenStudy (anonymous):

why did you leave me

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

it's ok theres the question

OpenStudy (anonymous):

5/5=1 75/5=15

OpenStudy (anonymous):

1/15

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

that it

OpenStudy (anonymous):

no but some of the questions are really long to type

OpenStudy (anonymous):

copy and paste

OpenStudy (anonymous):

i cant im on a different computer than what the questions are on

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

wait heres one

OpenStudy (anonymous):

write five and three eighths as an improper fraction and as a mixed number a. 43/8 and 5 3/8 b.48/8 and 3 5/8 c.37/8 and 5 3/8 d. 29/8 and 3 5/8

OpenStudy (anonymous):

43/8 5*8+3=43/8

OpenStudy (anonymous):

pasty needs 3 1/3 cups of flour for a recipe. write the number of cups needed as an improper fraction. a.11/3 b.9/3 c.10/3 d.7/3

OpenStudy (anonymous):

3*3=9+1=10 10/3

OpenStudy (anonymous):

list multiples to find the LCM of 27 and 45 a. 45 b.675 c.1,215 d.135

OpenStudy (anonymous):

the smallest number they both reach is 675

OpenStudy (anonymous):

write the decimal as a fraction or a mixed number in simplest form. 4.28 a.4 3/10 b.4 4/25 c.4 7/20 d.4 7/25

OpenStudy (anonymous):

im not good at transferring it back i cant do this one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

write an equivalent fraction for 6x + 5y + 7x a.18xy b.30xy c.11xy + 7x d.13x +5y

OpenStudy (anonymous):

d

OpenStudy (anonymous):

combine same letters

OpenStudy (anonymous):

write an equivalent expression for 8x + 6 using the distributive property. a.2(2x + 3) b.2(4x+3) c.8(x+6) d.x(8+6)

OpenStudy (anonymous):

b and how many questions u got left

OpenStudy (anonymous):

i wont give you any more

OpenStudy (anonymous):

can i get a medal

OpenStudy (anonymous):

how do i give you one

OpenStudy (anonymous):

click best responce

OpenStudy (anonymous):

can u find it @poopbrains